OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3)

OpenStudy (saifoo.khan):

4(x - 6) = -2(x - 3) divide 2 on both sides. 2(x-6) = -1 (x-3) 2x - 12 = -x + 3 2x + x = 12 + 3 3x = 15 x = 15/3 x = 5

OpenStudy (anonymous):

use the distributive property. a(b+c)=ab +ac

OpenStudy (anonymous):